derivatives order 0 Find higher the 2 la f(x) 3x + 72 tl Find f(x) b f(x) = x, (4) f (x) find f(x) = 205- 2 find f(5 (x) 2
Answer #1

(a) fons 3r?+73+1 fina d (332+72+1) an = 16] = 3(20) +701140 62+7 f62012 d tone) dr6u+7) ca dre Glinto of cus=r=1/2 = 6 f